The four tires of an automobile are inflated to a gauge pressure of 2.0x10^5 Pa. Each tire has an area of 0.024m^2 in contact with the ground. Determine the weight of the automobile.
Answer #1
so i do 2.0x10^5Pa x (0.024m^2)(4)??
answered by: daf
Answer #2
Weight=pressuretire*area (all four tires)
answered by: annonomus
Answer #3
would the answer be 19200?
answered by: Brandy
Answer #4
so i do 2.0x10^5Pa x (0.024m^2)(4)??
answered by: Meosha
Answer #5

P=F/A so PA=F. (2.0x10^5)(.024)=4800N 

F=ma (in this case a is acceleration due to gravity 9.8) F=mg 4800=9.8m 

m=489.8 kg 

4m=1959.2 kg